OpenStudy (anonymous):

How Many Solutions Does This Problem Have? –4y + 4 = x 2x + 8y = 8

OpenStudy (anonymous):

2, b/c it has 2 variables, with exponents of 1 on each variable.

OpenStudy (anonymous):

actually if u try to solve this it would be no solution, b/c one equation is a multiple of the other. (-4y+4=x)*2 = -8y+8=2x, which is the same as 2x+8y=8. Which coincidentally the same as equation 2

OpenStudy (anonymous):

Both equations represent the same so as we have just one equation with 2 variables there are infinite solutions, depends upon the value you give to x

OpenStudy (anonymous):

for example: x = 0 y = 1 x = 1 y = 3/4 etcetera

OpenStudy (anonymous):

well it was a pleasure to help you saludos desde Perú